I need to make 50 ml of a 2 percent dextrose solution from a 30 percent dextrose stock. How much diluent will I need ?

Answer by mananth(16946)

Dextrose
percent ---------------- quantity
Dilutant 0.00% ---------------- x ml
Dextrose 30.00% ------ 50 - x ml
Mixture 2.00% ---------------- 50
50
0.00% x + 30.00% ( 50 - x ) = 2.00% * 50
0 x + 30 ( 50 - x ) = 100
0 x + 1500 - 30 x = 100
0 x - 30 x = 100 - -1500
-30 x = -1400
/ -30
x = 46.67 ml 0.00% Dilutant
3.33 ml 30.00% Dextrose

m.ananth@hotmail.ca